My daugher and her friend are having a challenge solving the followin:
Johnstown High School is selling candy to raise money. M&M's sell for $1.50 per box and Reese's sell for $2 per box. Sheri sold a total of 50 boxes of candy. She collected $95.
Let x represent the number of boxes of M&M's
Let y represent the number of boxes of Reese's
Solve the system to determine how many of each type of candy did she sell?

Answer by Paul(988)

x+y=50
y=50-x

1.50x+2.00y=95 (subsitute the known value)
1.50x+2(50-x)=95
1.50x+100-2x=95
-0.5x=-5
x = 10
Hence, she sold 10 boxes of M&Ms and 40 boxes of Reese.